A boulder of mass 100 kg is pulled in some direction with a net force whose components are


A.

B.

C.

D.

= 100 N, 𝐹 = 200 N.


1
Expert's answer
2022-01-25T06:50:05-0500

Explanation & Calculations


  • Here it sounds like it asks for the net force.
  • Normally, the components are taken orthogonal to each other. "\small 100\,N\,\angle{90^\circ}\,200\,N".
  • Then, using the Pythagorean theorem, the resultant can be calculated as follows.

"\qquad\qquad
\begin{aligned}
\small R&=\small \sqrt{100^2+200^2}\\
&=\small 223.6\,N
\end{aligned}"

  • The angle created with the "\small 200\,N" can be calculated as "\small \tan\theta=\frac{100\,N}{200\,N}" .
  • However, it is impossible to provide a precise answer without seeing the 4 options.